OpenStudy (anonymous):

A cylinder 4in. In diameter and 6in height holds 1Ib of oatmeal. To the nearest ounce, how much oatmeal will a similar 10in high cylinder hold?

hartnn (hartnn):

Volume is directly proportional to height \((V= \pi r^2h)\) so, \(V \alpha h \implies \dfrac{V_2}{V_1} = \dfrac{h_2}{h_1}\) where, V1 = 1 , h1= 6 , h2= 10 just plug in and find V2